OpenStudy (anonymous):

You want to sell a certain number n of items in order to maximize your profit. Market research tells you that if you set the price at $1.50, you will be able to sell 5000 items, and for every 10 cents you lower the price below $1.50 you will be able to sell another 1000 items. Suppose that your fixed costs (“start-up costs”) total $2000, and the per item cost of production (“marginal cost”) is $0.50. Find the price to set per item and the number of items sold in order to maximize profit, and also determine the maximum profit you can get.

OpenStudy (anonymous):

this is what i got so far: r(x)=nx---> revenue for selling n items at x dollars c(x)=2000+0.5n ---> cost n(x)=5000 + 1000(1.5- 0.10x) these are my equations, i calculated it until the end, but i got wrong answer. i got x=1.55 or x=33.95 but the right answer is x=1.25

OpenStudy (anonymous):

let y be the price of product if n=5000 +1000x then y=1.5-0.1x and our profit (that we want to maximize) is: (5000+1000x)(1.5-0.1x-0.5)-2000 =(5000+1000x)(1-0.1x)-2000 =5000-500x+1000x-100x2-2000 =-100x2+500x+3000 (first derivative)->-200x+500=0 [second derivative: -200 (and we confirm that the function has maximum point)] ->x=-500/-200=2.5 price to set: 1.25 number of items: 7500 our profit: 3625